OpenStudy (anonymous):

a bottle holds 48 tsp of vanilla. the amount A of vanilla remaining in the bottle decreases by 2 tsp per batch B of cookies write a function rule to represent this situation. how much vanilla remains after 12 batches of cookies?

OpenStudy (anonymous):

f[B]=48-2B B=12 f[12]=48-2(12)

OpenStudy (anonymous):

=24